### v1

# A Polynomial-Time Algorithm for EFX Orientations of Chores

Abstract
This paper addresses the problem of finding EFX orientations of graphs of chores, in which each vertex corresponds to an agent, each edge corresponds to a chore, and a chore has zero marginal utility to an agent if its corresponding edge is not incident to the vertex corresponding to the agent. Recently, Zhou et al. (IJCAI, 2024) analyzed the complexity of deciding whether graphs containing a mixture of goods and chores admit EFX orientations, and conjectured that deciding whether graphs containing only chores admit EFX orientations is NP-complete. In this paper, we resolve this conjecture by exhibiting a polynomial-time algorithm that finds an EFX orientation of a graph containing only chores if one exists, even if the graph contains self-loops. Remarkably, our first result demonstrates a surprising separation between the case of goods and the case of chores, because deciding whether graphs containing only goods admit EFX orientations of goods was shown to be NP-complete by Christodoulou et al. (EC, 2023). In addition, we show the analogous decision problem for multigraphs to be NP-complete.
1 Introduction
Consider the problem of dividing a set of items among a group of independent agents with the stipulation that each item must be given to exactly one agent and cannot be shared. This problem can be used to model a wide variety of settings, such as assigning office space in a building to employees, allocating classrooms to multiple simultaneous lectures, and even divorce settlements between spouses. The classical example of dividing a single good between two agents shows that it is not always possible to find an envy-free allocation, necessitating the search for alternative criteria of fairness.
The most prominent such criterion is envy-freeness up to any good (EFX) introduced by Caragiannis et al. Caragiannis et al. (2019) and Gourvès et al. Gourvès et al. (2014). Essentially, because we cannot hope to achieve envy-freeness in general, we allow for an agent $i$ to envy another agent $j$, as long as the envy can be alleviated by ignoring any one good allocated to $j$. The existence problem of EFX allocations has proven to be notoriously difficult and has even been referred to as ”fair division’s most enigmatic question” by Procaccia Procaccia (2020). Many attempts have been made toward this problem by limiting the types of instances. For example, EFX allocations of goods exist if the agents have identical utility functions or if the agents have identical ranking of the goods Plaut and Roughgarden (2020), if the agents have lexicographic utility functions Hosseini et al. (2023), or if the number of agents is at most 3 Chaudhury et al. (2020); Akrami et al. (2022).
Another possibility is to limit the number of agents who are interested in each item. The assumption that each item provides non-zero marginal utility to at most two agents leads to the graphical instances introduced by Christodoulou et al. Christodoulou et al. (2023). These instances can be represented by a graph $G$ in the following way. Each agent is represented by a vertex, and each item is represented by an edge between the two agents for whom it possibly provides non-zero marginal utility. Thus, whenever an edge is not incident to a vertex, it provides zero marginal utility to the agent corresponding to that vertex. Because instances are represented as graphs, it is natural to ask whether EFX orientations exist, that is, whether it is possible to allocate each edge to one of its endpoints while satisfying the EFX condition. Unfortunately, it is NP-complete to decide whether EFX orientations exist for instances containing only goods Christodoulou et al. (2023).
The recent work due to Zhou et al. Zhou et al. (2024) explored the above problem in further detail, studying the complexity of determining whether EFX orientations exist when both goods and chores are present (i. e. the mixed manna setting). Under this setting, the original definition of EFX can be adapted in four different ways depending on whether a good is ignored in the envied agent’s bundle (similar to the original definition of EFX) or a chore is ignored in the envious agent’s bundle, and also depending on if the ignored item must have non-zero marginal utility. Zhou et al. Zhou et al. (2024) analyzed the complexity of deciding whether EFX orientations and allocations exist for each of the four variants of EFX.
Their paper concludes with a brief discussion on the special case of instances containing only chores. Two different variants of EFX can be defined for such instances. Specifically, let $\pi$ be an allocation, $\pi_{i}$ denote the bundle of agent $i$, and $u_{i}$ denote the utility function of agent $i$. We say that $\pi$ is
  • EFX_ if for each pair of agents $i\neq j$ such that $i$ envies $j$ and each $e\in\pi_{i}$ such that $u_{i}(e)<0$, we have $u_{i}(\pi_{i}\setminus\{e\})\geq u_{i}(\pi_{j})$; and
  • EFX0 if for each pair of agents $i\neq j$ such that $i$ envies $j$ and each $e\in\pi_{i}$ such that $u_{i}(e)\leq 0$, we have $u_{i}(\pi_{i}\setminus\{e\})\geq u_{i}(\pi_{j})$.
The difference is that while EFX_ only requires the envy be alleviated when ignoring chores with strictly negative marginal utility, EFX0 requires that the envy be alleviated even if we ignore a chore with zero marginal utility. Clearly, EFX0 is a strictly stronger condition than EFX_.
Zhou et al. Zhou et al. (2024) found a simple polynomial-time algorithm that decides whether a graph of chores admits an EFX_ orientation. On the other hand, they conjectured that deciding whether a graph of chores admits an EFX0 orientation is NP-complete, similar to the case involving only goods.
1.1 Our Contribution
In this paper, we resolve the conjecture made by Zhou et al. Zhou et al. (2024) by giving a polynomial-time algorithm that decides the existence of an EFX0 orientation of chores and outputs one if it exists. Our algorithm is also able to handle graphs with self-loops and is essentially based on reducing the problem of finding EFX0 orientations to 2SAT, which can be solved in polynomial time Aspvall et al. (1979). Remarkably, our result stands in contrast with the case of goods, which is NP-complete Christodoulou et al. (2023).
Theorem 1.
There exists a $O(|V(G)|^{2})$-time algorithm that decides whether a graph $G$ of chores admits an EFX0 orientation and outputs one if it exists.
In contrast, we show that the analogous problem for multigraphs is NP-complete.
Theorem 2.
Deciding whether a multigraph $G$ of chores admits an EFX0 orientation is NP-complete, even if $G$ contains only two vertices and agents have additive utility functions.
1.2 Organization
We introduce the model in Section 2. Our algorithm for finding EFX0 orientations of graphs is given in Section 3 (Theorem 1). Finally, we prove our hardness result for multigraphs in Section 4 (Theorem 2).
2 Preliminaries
2.1 The EFX0 Orientation Problem
An instance of EFX0-Orientation is an ordered pair $(G,u)$ where $G=(V(G),E(G))$ is a graph and $u$ is a collection of $|V(G)|$ utility functions $u_{i}:\mathcal{P}(E(G))\to\mathbb{R}_{\leq 0}$. The graph $G$ possibly contains self-loops. Each vertex corresponds to an agent and each edge corresponds to a chore. For each agent $i$, the utility function $u_{i}$ represents its preferences by assigning a numerical value to each subset of the edges. In the case of a single edge $e$, we write $u_{i}(e)$ in place of $u_{i}(\{e\})$ for brevity. We further assume that whenever a vertex $i$ is not incident to an edge $e$, the edge $e$ has zero marginal utility to $i$, i. e. $u_{i}(S\cup\{e\})=u_{i}(S)$ for any set $S\subseteq E(G)$. In addition, we make the following two minimal assumptions on each utility function $u_{i}$:
  • $u_{i}$ is monotone, i. e., $u_{i}(S)\leq u_{i}(T)$ whenever $T\subseteq S$.
  • If $u_{i}(e)=0$ for each $e\in S$, then $u_{i}(S)=0$.
Importantly, our assumptions on utility functions are strictly more general than additivity.
An allocation $\pi=(\pi_{1},\pi_{2},\dots,\pi_{|V(G)|})$ is an ordered collection of $|V(G)|$ pairwise disjoint subsets of $E(G)$ whose union is exactly $E(G)$. For each $i$, the subset $\pi_{i}$ is called the bundle of agent $i$. We are interested in a special kind of allocations called orientations, which are allocations that can be represented as a graph orientation $O$ of $G$ such that $e\in\pi_{i}$ if and only if $e$ is directed toward $i$ in $O$. We say an agent $i$ strongly envies an agent $j$ if there exists an edge $e\in\pi_{i}$ such that $u_{i}(\pi_{i}\setminus\{e\})<u_{i}(\pi_{j})$. An allocation is said to be envy-free up to any edge (chore) (EFX0) if no agent strongly envies another. In other words, any envy $i$ has for $j$ can be alleviated by ignoring any edge $e$ in the bundle of agent $i$.
Given an instance $(G,u)$ of the above form, EFX0-Orientation asks whether $G$ admits an EFX0 orientation.
We are particularly interested in a special case of EFX0-Orientation in which every edge has a certain property. An edge $e=\{i,j\}$ in a graph is said to be objective if $u_{i}(e)=0$ if and only if $u_{j}(e)=0$. In other words, either $e$ has zero marginal utility for both of its endpoints, or $e$ has strictly negative marginal utility for both of them. If every edge is objective, we say $(G,u)$ is an objective instance of EFX0-Orientation. When considering such instances, we find it convenient to refer to edges with zero (resp. negative) marginal utility for both endpoints as dummy edges (resp. negative edges). In addition, we define a negative component of $G$ to be a maximal vertex-induced subgraph $K$ such that for any two vertices $i,j$ of $K$, there exists a path in $K$ between $i$ and $j$ that only contains negative edges. The negative components of $G$ form a partition of $V(G)$.
2.2 Related Decision Problems
We define some relevant decision problems. Let $H$ be a graph, $P=\{P_{1},P_{2},\dots,P_{k}\}$ be a set of pairwise disjoint subsets of $V(H)$, and $D\subseteq V(H)$ be a subset of vertices. The set $P$ is not necessarily a partition because not every vertex needs to be in $P$. We say a subset $C\subseteq V(H)$ is a $(P,D)$-vertex cover of $H$ if:
  • Each edge of $H$ is incident to at least one vertex in $C$;
  • $|C\cap P_{\ell}|\leq 1$ for each $P_{\ell}\in P$; and
  • $C\cap D=\emptyset$.
Essentially, $C$ is a vertex cover that contains at most one vertex from each $P_{i}$ and no vertex in $D$.
Definition 1 (PD-Vertex-Cover).
Given a tuple $(H,P,D)$ where $H$ is a graph, $P=\{P_{1},P_{2},\dots,P_{k}\}$ is a set of pairwise disjoint subsets of $V(H)$, and $D\subseteq V(H)$, PD-Vertex-Cover asks whether there exists a $(P,D)$-vertex cover $C$ of $H$.
Definition 2 (2SAT).
The 2SAT problem is a special case of SAT in which each clause consists of one or two literals.
There exists an algorithm that decides 2SAT and outputs a satisfying assignment if one exists, with running time linear in the number of variables and clauses Aspvall et al. (1979).
Definition 3 (Partition).
Given a set $S=\{s_{1},s_{2},\dots,s_{k}\}$ of natural numbers, Partition asks whether there exists a partition $S=S_{a}\cup S_{b}$ such that $\sum_{x\in S_{a}}x=\sum_{x\in S_{b}}x$.
3 EFX0 Orientations of Graphs
In this section, we present our polynomial-time algorithm for EFX0-Orientation. Our algorithm is essentially the result of a reduction from EFX0-Orientation to 2SAT, and is composed of three nested routines, the outermost of which is the main algorithm FindEFXOrientation.
To avoid forward referencing in the following subsections, we present the proofs of correctness of the algorithms in their nested order, starting with most deeply nested one.
3.1 The Subroutine FindPDVertexCover
The subroutine FindPDVertexCover (Algorithm 1) accepts an instance $(H,P,D)$ of PD-Vertex-Cover as input, and outputs a $(P,D)$-vertex cover $C$ of $H$ if one exists and false otherwise.
Recall that in an instance $(H,P,D)$ of PD-Vertex-Cover, $H$ is a graph, $P=\{P_{1},P_{2},\dots,P_{k}\}$ is a set of pairwise disjoint subsets of $V(H)$, and $D\subseteq V(H)$. The algorithm starts by constructing an instance $\phi$ of 2SAT on Lines 110, which contains a Boolean variable $x_{i}$ for each vertex $i$ of $H$ and three types of clauses:
  • (type 1) the clause $\{x_{i},x_{j}\}$ for each edge $e=\{i,j\}$ of $H$ (possibly $i=j$),
  • (type 2) the clause $\{\overline{x_{i}},\overline{x_{j}}\}$ for each set $P_{\ell}\in P$ and each pair of distinct vertices $i,j\in P_{\ell}$, and
  • (type 3) the clause $\{\overline{x_{i}}\}$ for each vertex $i\in D$.
Then, the algorithm determines whether $\phi$ admits a satisfying assignment $f$ on Line 11. If so, the algorithm returns the set $C\coloneqq\{i\in V(H):f(x_{i})=\texttt{true}\}$. Otherwise, the algorithm returns false.
[…]
Input: An instance $(H,P,D)$ of PD-Vertex-Cover.
Output: A $(P,D)$-vertex cover $C$ of $H$ if one exists and false otherwise.
1:$\phi\leftarrow\emptyset$
2:for each vertex $i$ of $H$ do
3:     create the Boolean variable $x_{i}$
4:for edge $e=\{i,j\}$ of $H$ do
5:     $\phi\leftarrow\phi\cup\{\{x_{i},x_{j}\}\}$triangleright type 1 clauses
6:for each set $P_{\ell}\in P$ do
7:     for each pair of distinct vertices $i,j\in P_{\ell}$ do
8:         $\phi\leftarrow\phi\cup\{\{\overline{x_{i}},\overline{x_{j}}\}\}$triangleright type 2 clauses
9:for each vertex $i\in D$ do
10:     $\phi\leftarrow\phi\cup\{\{\overline{x_{i}}\}\}$triangleright type 3 clauses
11:$f\leftarrow\textsc{2SAT}(\phi)$
12:if $f=\texttt{false}$ then $\triangleright$ $\phi$ has no satisfying assignment
13:     return false
14:return $C\coloneqq\{i\in V(H):f(x_{i})=\texttt{true}\}$
Lemma 1.
The following statements hold for FindPDVertexCover (Algorithm 1):
  1. 1.
    If $H$ admits a $(P,D)$-vertex cover, then $\phi$ admits a satisfying assignment; and
  2. 2.
    If $\phi$ admits a satisfying assignment $f$, then $C\coloneqq\{i\in H:f(x_{i})={\texttt{true}}\}$ is a $(P,D)$-vertex cover of $H$. Otherwise, FindPDVertexCover outputs false.
  3. 3.
    FindPDVertexCover runs in $O(|V(H)|^{2})$ time.
Proof.
(1:) Suppose $H$ admits a $(P,D)$-vertex cover $C$. We claim that the truth assignment $f$ in which $f(x_{i})=\texttt{true}$ if and only if $i\in C$ is a satisfying assignment for $\phi$. Since $C$ is a $(P,D)$-vertex cover of $H$, each edge $e$ of $H$ is incident to at least one vertex in $C$. In other words, for each edge $e=\{i,j\}$ of $H$, we have $i\in C$ or $j\in C$. Hence, at least one of $x_{i},x_{j}$ is true under $f$, so the type 1 clause $\{x_{i},x_{j}\}$ is satisfied.
On the other hand, the definition of $(P,D)$-vertex covers ensures $|C\cap P_{\ell}|\leq 1$ for each $P_{\ell}\in P$. Hence, each $P_{\ell}$ contains at most one vertex $i$ such that $f(x_{i})=\texttt{true}$, so all type 2 clauses are satisfied.
Finally, because $C$ is a $(P,D)$-vertex cover, we have $C\cap D=\emptyset$. Hence, for each $i\in D$, we have $i\notin C$, so $f(x_{i})=\texttt{false}$, ensuring all type 3 clauses are satisfied.
(2:) Suppose $\phi$ admits a satisfying assignment $f$. We claim that $C\coloneqq\{i\in V(H):f(x_{i})=\texttt{true}\}$ is a $(P,D)$-vertex cover of $H$. Since the type 1 clause $\{x_{i},x_{j}\}$ is satisfied under $f$, the edge $e=\{i,j\}$ of $H$ is incident to at least one vertex of $C$, namely, $i$ or $j$. Since the type 2 clauses are satisfied, for each fixed $P_{\ell}\in P$, there is at most one vertex $i\in P_{\ell}$ such that $x_{i}$ is true under $f$. Hence, $|C\cap P_{\ell}|\leq 1$. Finally, the type 3 clauses ensure that for each vertex $i\in D$, we have $f(x_{i})=\texttt{false}$, so $i\notin C$. Thus, $C\cap D=\emptyset$.
Otherwise, $\phi$ admits no satisfying assignment and FindPDVertexCover outputs false on Line 13.
(3:) Constructing $\phi$ on Lines 110 takes $O(|V(H)|^{2})$ time because it contains $|V(H)|$ variables and $O(|V(H)|^{2})$ clauses. Verifying whether $\phi$ admits a satisfying assignment on Line 11 takes time linear in the number of variables and clauses by using the algorithm due to Aspvall et al. Aspvall et al. (1979), which is $O(|V(H)|^{2})$. Finally, constructing $C$ takes $O(|V(H)|)$ time by checking each vertex of $H$. Thus, FindPDVertexCover runs in $O(|V(H)|^{2})$ time. ∎
3.2 The Subroutine FindEFXOrientObj
The subroutine FindEFXOrientObj (Algorithm 2) takes as input an objective instance $(G^{o},u^{o})$ of EFX0-Orientation and outputs an EFX0 orientation of $G^{o}$ if one exists and false otherwise. Recall that in an objective instance, all edges are objective, and we distinguish dummy edges that provide zero marginal utility to both endpoints and negative edges that provide negative marginal utility to both endpoints. Recall also that a negative component of $G^{o}$ is a maximal vertex-induced subgraph $K$ such that for any two vertices $i,j$ of $K$, there exists a path in $K$ between $i$ and $j$ that only contains negative edges.
FindEFXOrientObj begins by using breadth-first search (BFS) to find the set $\mathcal{K}$ of negative components of $G^{o}$ on Lines 17. It then constructs the instance $(H,P,D)$ of PD-Vertex-Cover on Lines 817 such that
  • $H$ is the graph with the same vertex set as $G^{o}$ and the set of dummy edges of $G^{o}$;
  • $P=\{K\in\mathcal{K}:K\text{ contains exactly }|K|-1\text{ negative edges}\}$; and
  • $D$ is the set of vertices that belong to a negative component $K$ of $G^{o}$ containing exactly $|K|$ negative edges.
In the case that $G^{o}$ contains a negative component $K$ with more than $|K|$ negative edges, FindEFXOrientObj returns false. Otherwise, each negative component $K$ contains exactly $|K|-1$ or $|K|$ negative edges. Via a call to FindPDVertexCover (Algorithm 1) on Line 18, the algorithm attempts to find a $(P,D)$-vertex cover $C$ of $H$. If unsuccessful, it returns false. Otherwise, it uses $C$ to construct an orientation $\pi^{o}$ of $G^{o}$ as follows.
First, for each dummy edge $e$ that is incident to a vertex $i$ of $C$, the algorithm allocates $e$ to $i$, with arbitrary tie-breaking when $e$ is incident to two vertices of $C$ (Lines 2223). Then, the algorithm orients the negative edges of $G^{o}$, one negative component $K$ at a time as follows. If $K$ has exactly $|K|$ negative edges, then the algorithm orients them on Line 26 in a way that each vertex of $K$ receives exactly one negative edge. This is possible because the $|K|$ negative edges of $K$ take the form of a spanning tree of $K$ with an extra edge, so there is a unique cycle $C$ of negative edges of $K$. By orienting $C$ cyclically, any remaining unoriented negative edges of $K$ form trees branching from $C$, so we can orient them in the direction away from $C$, ensuring each vertex of $K$ receives exactly one negative edge.
Otherwise, $K$ has exactly $|K|-1$ negative edges. The algorithm proceeds to Lines 2932 and chooses a vertex $r$ of $K$ to be one that has received a dummy edge in the previous step if such a vertex exists, or any arbitrary vertex of $K$ if no vertex of $K$ has received a dummy edge. Once $r$ is chosen, the algorithm orients the $|K|-1$ negative edges of $K$ in the direction away from $r$. This causes these negative edges to form a spanning rooted tree of $K$ rooted at $r$, ensuring that each vertex of $K$ receives at most one negative edge. FindEFXOrientObj returns the resulting orientation $\pi^{o}$ of $G^{o}$.
[…]
Input: An objective instance $(G^{o},u^{o})$ of EFX0-Orientation.
Output: An EFX0 orientation $\pi^{o}$ of $G^{o}$ if one exists and false otherwise.
1:$\mathcal{K}\leftarrow\emptyset$
2:unvisited $\leftarrow V(G^{o})$
[…]
4:     $v\leftarrow$ any vertex in unvisited
5:     $K\leftarrow$ the set of vertices of $G^{o}$ reachable via BFS from $v$, when only negative edges are considered.
6:     $\mathcal{K}\leftarrow\mathcal{K}\cup\{K\}$
7:     unvisited $\leftarrow$ unvisited $\setminus K$
8:$H\leftarrow$ graph with vertex set $V(G^{o})$ and edge set $\{e\in E(G^{o}):e\text{ is a dummy edge}\}$
9:$P\leftarrow\emptyset$
10:$D\leftarrow\emptyset$
11:for each negative component $K$ of $G^{o}$ do
12:     if $K$ has $>|K|$ negative edges then
13:         return false
14:     else if $K$ has $|K|$ negative edges then
15:         $D\leftarrow D\cup V(K)$
16:     $\triangleright$ $K$ has $|K|-1$ negative edges
17:         $P\leftarrow P\cup\{V(K)\}$
18:$C\leftarrow\textsc{FindPDVertexCover}(H,P,D)$triangleright Algorithm 1
19:if $C=\texttt{false}$ then
20:     return false
21:$\pi^{o}\leftarrow(\pi_{1}^{o},\pi_{2}^{o},\dots,\pi^{o}_{|V(G^{o})|})$ where each $\pi^{o}_{i}=\emptyset$
22:for $i\in C$ do
23:     $\pi^{o}_{i}\leftarrow$ all unallocated dummy edges of $G^{o}$ incident to $i$
24:for each negative component $K$ of $G^{o}$ do
25:     if $K$ has exactly $|K|$ negative edges then
26:         Let $\pi^{o}$ orient them so that each vertex of $K$ receives exactly one negative edge.
27:     $\triangleright$ $K$ has exactly $|K|-1$ negative edges
28:         if $\exists i\in K$ such that $\pi_{i}^{o}$ contains a dummy edge then
29:              $r\leftarrow$ such a vertex $i\in K$
30:         else
31:              $r\leftarrow$ any arbitrary vertex $i\in K$
32:         Let $\pi^{o}$ orient the negative edges of $K$ as the directed tree where all edges point away from $r$.
33:return $\pi^{o}$
Proposition 1.
Let $(G^{o},u^{o})$ be an objective instance of EFX0-Orientation. An orientation $\pi^{o}$ of $G^{o}$ is EFX0 if and only for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge.
Proof.
Let $\pi^{o}=(\pi_{1}^{o},\pi_{2}^{o},\dots,\pi_{|V(G^{o})|})$ be an orientation of $G^{o}$ and suppose that for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. If $\pi^{o}$ contains a unique edge $e$ directed toward $i$, then we have $u_{i}^{o}(\pi^{o}_{i}\setminus\{e\})=u_{i}^{o}(\emptyset)=0\geq u_{i}^{o}(\pi_{j}^{o})$ for any agent $j$. Otherwise, every edge directed toward $i$ is a dummy edge, so $u_{i}^{o}(\pi_{i}^{o})=0\geq u_{i}^{o}(\pi_{j}^{o})$ for any $j$. Hence, no agent strongly envies another and $\pi^{o}$ is EFX0.
Conversely, suppose that for some vertex $i$, the orientation $\pi^{o}$ contains two edges $e,e^{\prime}$ directed toward $i$ and not every edge directed toward $i$ is a dummy edge. Without loss of generality, assume $u_{i}^{o}(e^{\prime})<0$. Let $j,j^{\prime}$ denote the other endpoints of the edges $e,e^{\prime}$, respectively. Note that $j\neq j^{\prime}$ because $G$ does not contain parallel edges. Also, $u_{i}^{o}(\pi_{j}^{o})=0$ because agent $i$ receives the edge $e$ between $i$ and $j$. Since $u_{i}^{o}(e^{\prime})<0$ and $\{e^{\prime}\}\subseteq\pi_{i}^{o}\setminus\{e\}$, the monotonicity of $u_{i}^{o}$ implies $u_{i}^{o}(\pi_{i}^{o}\setminus\{e\})\leq u_{i}^{o}(e^{\prime})<0=u_{i}^{o}(\pi_{j}^{o})$. It follows that $\pi^{o}$ is not EFX0. ∎
Lemma 2.
The following statements hold for FindEFXOrientObj (Algorithm 2):
  1. 1.
    If $G^{o}$ admits an EFX0 orientation, then no negative component $K$ of $G^{o}$ contains $>|K|$ negative edges and $H$ admits a $(P,D)$-vertex cover.
  2. 2.
    If $C$ is a $(P,D)$-vertex cover of $H$, then $\pi^{o}$ is an EFX0 orientation of $G^{o}$.
  3. 3.
    FindEFXOrientObj runs in $O(|V(G^{o})|^{2})$ time.
Proof.
(1:) Suppose $G^{o}$ admits an EFX0 orientation $\pi^{o}$. First, we show that no negative component $K$ of $G^{o}$ contains $>|K|$ negative edges. Fix a negative component $K$. By Proposition 1, for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. In particular, this is true for each vertex in $K$, so $\pi^{o}$ contains at most $|K|$ negative edges directed toward a vertex of $K$. On the other hand, every negative edge of $K$ is directed toward a vertex of $K$ because $\pi^{o}$ is an orientation. It follows that $K$ contains at most $|K|$ negative edges.
We claim that the set $C$ of vertices $i$ of $G^{o}$ such that $\pi^{o}$ contains a dummy edge directed toward $i$ is a $(P,D)$-vertex cover of $H$. First we show that each edge of $H$ is incident to at least one vertex in $C$. Let $e$ be an edge of $H$. By the definition of $H$, the edge $e$ is a dummy edge of $G^{o}$. Let $i$ be the endpoint toward which $e$ is directed in the orientation $\pi^{o}$ of $G^{o}$. The definition of $C$ implies $i\in C$, so $e$ is incident to at least one vertex in $C$.
Next, we show that $|C\cap K_{\ell}|\leq 1$ for each $K_{\ell}\in P$. Fix any negative component $K_{\ell}$ of $G^{o}$. Since $K_{\ell}$ is a negative component, it contains at least $|K_{\ell}|-1$ negative edges. Since $\pi^{o}$ is EFX0, Proposition 1 implies that for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. Since $K_{\ell}$ contains at least $|K_{\ell}|-1$ negative edges and each vertex of $K$ can have at most one negative edge directed towards it, the pigeonhole principle implies that $K_{\ell}$ contains at most one vertex to which dummy edges are directed in $\pi^{o}$. In other words, at most one vertex of $K_{\ell}$ is in $C$, i. e., $|C\cap K_{\ell}|\leq 1$.
Finally, we show that $C\cap D=\emptyset$. Suppose $i\in D$. Because $i\in D$, the negative component $K$ to which $i$ belongs contains $|K|$ negative edges. Proposition 1 implies that for each vertex $j$, the orientation $\pi^{o}$ contains a unique edge directed toward $j$ or every edge directed toward $j$ is a dummy edge. In particular, each vertex of $K$ has at most one negative edge directed toward it. By the pigeonhole principle, each vertex of $K$ has exactly one negative edge of $K$ directed toward it. In particular, this is true for $i$, so Proposition 1 implies no dummy edge is directed toward $i$. Thus, $i\notin C$. It follows that $C\cap D=\emptyset$ and $H$ indeed admits a $(P,D)$-vertex cover.
(2:) Suppose the algorithm found a $(P,D)$-vertex cover $C$ of $H$ on Line 18. First, we show that $\pi^{o}$ is indeed orients every edge of $G^{o}$. Recall that $H$ is the graph with the same vertex set as $G^{o}$ and the set of dummy edges of $G^{o}$. Since $C$ is a $(P,D)$-vertex cover of $H$, each edge of $H$ (thereby each dummy edge of $G^{o}$) is incident to at least one vertex in $C$. It follows that every dummy edge of $G^{o}$ is allocated on Lines 2223. On the other hand, let $e$ be a negative edge of $G^{o}$ and $K$ be the negative component containing $e$. Since the algorithm has successfully found the $(P,D)$-vertex cover $C$ of $H$ on Line 18, it did not terminate on Line 13. So, $K$ contains at most $|K|$ negative edges and the algorithm orients every negative edge on either Line 26 or 32. Thus, $\pi^{o}$ indeed orients every edge of $G^{o}$.
It remains to show that $\pi^{o}$ is EFX0. We claim that each vertex $i$ that receives a dummy edge on Line 23 receives no negative edges in the final for-loop (Line 2432). Let $C^{\prime}\subseteq C$ be the set of vertices that receive a dummy edge on Line 23 and consider any $i\in C^{\prime}$. Since $C$ is a $(P,D)$-vertex cover of $H$, we have $C\cap D=\emptyset$, so $i\neq D$. Because $D$ is the set of vertices that belong to a negative component $K$ of $G^{o}$ containing exactly $|K|$ negative edges, $i$ belongs to a negative component $K$ of $G^{o}$ containing exactly $|K|-1$ negative edges. Moreover, because $C$ is a $(P,D)$-vertex cover, we have $|P_{\ell}\cap C|\leq 1$ for each $P_{\ell}\in P$, so $|K\cap C|\leq 1$ because $K\in P$. Hence, $i$ is the unique vertex in $K\cap C$, implying that $i$ is the only vertex of $K$ that receives a dummy edge on Line 23. Thus, during the iteration of the final for-loop corresponding to the negative component $K$ containing $i$, the vertex $i$ is the unique choice for $r$, so the algorithm sets $r=i$ and orients the negative edges of $K$ as the directed tree where all edges point away from $i$, ensuring that $i$ receives no negative edges in the final for-loop, proving our claim.
On the other hand, it is easy to see that each vertex clearly receives at most one negative edge in the final for-loop. In particular, each vertex $i\notin C^{\prime}$ receives no dummy edges and at most one negative edge in the final for-loop.
Consequently, for each vertex $i$, either $i\in C^{\prime}$ and $i$ receives no negative edges in the final for-loop, or $i\notin C^{\prime}$ and $i$ receives at most one negative edge in the final for-loop. In either case, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. By Proposition 1, $\pi^{o}$ is EFX0.
(3:) Finding the set $\mathcal{K}$ of negative components of $G^{o}$ on Lines 17 can be done using a BFS using $O(|V(G^{o})|+|E(G^{o})|)$ time. Constructing $(H,P,D)$ on Lines 817 requires iterating through each negative component $K$ of $G^{o}$ and counting the number of pairs of vertices of $K$ with a negative edge in between, taking $O(|E(G^{o})|)$ time. By Lemma 1(3), FindPDVertexCover on Line 18 runs in $O(|V(H)|^{2})$ time. The second for-loop (Lines 2223) takes $O(|V(H)|^{2})$ time because it checks each vertex of $C\subseteq V(H)$ and its incident edges.
Finally, consider the last for-loop (Lines 2432). This for-loop iterates at most $|V(G^{o})|$ times, each time orienting the negative edges within a negative component $K$ of $G^{o}$. We claim it can be implemented using at most two BFSs for each $K$. Suppose $K$ has $|K|$ negative edges. The negative edges of $K$ take the form of a spanning tree of $K$ together with an extra edge, i. e., they form a unique cycle $C_{K}$ and the set of negative edges not on $C_{K}$ induce a forest. In order to orient them in a way so that each vertex of $K$ receives exactly one negative edge, we can first use BFS to find the cycle $C_{K}$ and orient it cyclically. Then, by identifying the vertices of $C_{K}$ together as a single supervertex and doing a BFS starting at this supervertex, we can orient the remaining negative edges of $K$ away from the cycle $C_{K}$. This ensures that each vertex of $K$ on $C_{K}$ receives exactly one negative edge of $C_{K}$, and each vertex of $K$ not on $C_{K}$ receives at most one negative edge. Together, this takes two BFSs on $K$.
Otherwise $K$ has only $|K|-1$ negative edges. Orienting them as a directed rooted tree can be done using a single BFS on $K$. This proves our claim. Thus, the final for-loop takes $O(|V(G^{o})|+|E(G^{o})|)$ time. By summing the above together and observing that $|V(H)|=|V(G^{o})|$, we see that FindEFXOrientObj runs in $O(|V(G^{o})|^{2})$ time. ∎
Lemma 3.
Given an objective instance $(G^{o},u^{o})$ of EFX0-Orientation, FindEFXOrientObj outputs an EFX0 orientation $\pi^{o}$ of $G^{o}$ if one exists and false otherwise in $O(|V(G^{o})|^{2})$ time.
Proof.
If $G^{o}$ admits an EFX0 orientation, then no negative component $K$ of $G^{o}$ contains $>|K|$ negative edges and $H$ admits a $(P,D)$-vertex cover. So, Line 18 finds a $(P,D)$-vertex cover $C$ of $H$ by Lemma 1 and FindEFXOrientObj outputs the EFX0 orientation $\pi^{o}$. Otherwise, $G^{o}$ does not admit an EFX0 orientation, so some negative component $K$ of $G^{o}$ contains $>|K|$ negative edges FindEFXOrientObj outputs false on Line 13, or $H$ does not admit a $(P,D)$-vertex cover and FindEFXOrientObj outputs false on Line 20. ∎
3.3 The Main Algorithm FindEFXOrientation
FindEFXOrientation accepts an instance $(G,u)$ of EFX0-Orientation as input and outputs an EFX0 orientation of $G$ if one exists and false otherwise.
It first constructs an objective instance $(G^{o},u^{o})$ of EFX0-Orientation by subdividing each non-objective edge of $G$ in the for-loop on Lines 415. Specifically, for each non-objective edge $e_{ij}=\{i,j\}$ of $G$, it first possibly exchanges the labels between $i$ and $j$ so as to ensure $u_{i}(e_{ij})=0$. It then subdivides $e_{ij}$ by creating a new vertex $k$ and two new edges $e_{ik}=\{i,k\},e_{jk}=\{j,k\}$, while defining the utilities of the new edges in $G^{o}$ as shown in Figure 1, in which $\beta\coloneqq u_{j}(e_{ij})$. While iterating through the edges of $G$, the algorithm also constructs the set $O$ of objective edges that were not subdivided, and the set $S$ of ordered pairs $(e_{ij},e_{ik})$ where $e_{ij}$ is subdivided edge of $G$ and $e_{ik}$ is the resulting edge of $G^{o}$ incident to $i$.
The algorithm then calls FindEFXOrientObj as a subroutine on Line 16, which attempts to find an EFX0 orientation of $G^{o}$. If none exists, the algorithm returns false on Line 18. Otherwise, the subroutine successfully finds an EFX0 orientation $\pi^{o}$ of $G^{o}$. Using $\pi^{o}$, the algorithm constructs an orientation $\pi$ of $G$ on Lines 1926 as follows. First, any edge $e_{ij}$ of $G$ that was not subdivided in the above step is an edge in both $G$ and $G^{o}$. We let $\pi$ orient every such edge in $G$ in the same way as $\pi^{o}$ orients them in $G^{o}$.
Otherwise, the edge $e_{ij}$ of $G$ was subdivided in the above step and corresponds to two fake edges $e_{ik},e_{jk}$ of $G^{o}$. Recall that by possibly exchanging the labels of $i$ and $j$ in the subdivision of $e_{ij}$ above (Line 7), we may assume $u_{i}^{o}(e_{ik})=u_{i}(e_{ij})=0$. We let $\pi$ orient $e_{ij}$ toward $i$ in $G$ if and only if $\pi^{o}$ orients $e_{ik}$ toward $i$ in $G^{o}$. At the end, the algorithm outputs the resulting orientation $\pi$ of $G$.
[…]
(a) the edge $e_{ij}$ in $G$
[…]
(b) after subdivision in $G^{o}$
Figure 1: Subdivision of $e_{ij}$ during the construction of $G^{o}$ by FindEFXOrientation. We write $\beta\coloneqq u_{j}(e_{ij})$ for clarity.
Lemma 4.
The following statements hold for FindEFXOrientation (Algorithm 3):
  1. 1.
    If $G$ admits an EFX0 orientation, then so does $G^{o}$; and
  2. 2.
    If $\pi^{o}$ is an EFX0 orientation of $G^{o}$, then $\pi$ is an EFX0 orientation of $G$.
Proof.
(1:) Suppose $\pi$ is an EFX0 orientation of $G$. We construct the orientation $\pi^{o}$ of $G^{o}$ as follows. For each edge $e$ of $G$ that was not subdivided during the construction of $G^{o}$, we let $\pi^{o}$ orient it in $G^{o}$ in the same direction that $\pi$ orients it in $G$. On the other hand, suppose $e_{ik}$ and $e_{jk}$ are a pair of edges of $G^{o}$ that resulted from the subdivision of the edge $e_{ij}$ of $G$. If $\pi$ orients $e_{ij}$ leftward in terms of Figure 1 (resp. rightward), then $\pi^{o}$ orients both $e_{ik}$ and $e_{jk}$ leftward (resp. rightward).
We claim $\pi^{o}$ is EFX0. For convenience, we refer to a vertex that is shared between both $G$ and $G^{o}$ as a real vertex, and a vertex that exists only in $G^{o}$ as a fake vertex. Fake vertices result from the subdivisions during the construction of $G^{o}$. It is easy to see that from the local perspective of each real vertex, the orientations $\pi$ and $\pi^{o}$ appear the same. Since no real vertex strongly envies a neighbour in $\pi$, no real vertex strongly envies a neighbour in $\pi^{o}$. On the other hand, the construction of $\pi^{o}$ ensures each fake vertex receives exactly one incident edge, so no fake vertex strongly envies a neighbour in $\pi^{o}$. Thus, $\pi^{o}$ is EFX0 as claimed.
(2:) Suppose $\pi^{o}$ is an EFX0 orientation of $G^{o}$. Proposition 1 implies for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. In particular, for any fake vertex $k$ in $G^{o}$ and the two edges $e_{ik},e_{jk}$ incident to it, $e_{ik}$ is a dummy edge and $e_{jk}$ is a negative edge, so at most one of them is directed toward $k$ in $\pi^{o}$. Hence, there are only three possibilities for how pairs of fake edges are oriented in $\pi^{o}$ (see Figure 2).
Consider the first two cases depicted in Figure 2. In Figure 2(a), both $e_{ik}$ and $e_{jk}$ are directed rightward in $\pi^{o}$, and $\pi$ orients the corresponding edge $e_{ij}$ rightward. In Figure 2(b), both $e_{ik}$ and $e_{jk}$ are directed leftward in $\pi^{o}$, and $\pi$ orients $\pi_{ij}$ leftward. In either case, the real vertices $i$ and $j$ see no difference between $\pi$ and $\pi^{o}$ locally in terms of the utilities they derive from the edges $e_{ij},e_{ik},e_{jk}$ between $i$ and $j$ that they receive, because $u_{i}(e_{ij})=u_{i}(e_{ik})$ and $u_{j}(e_{ij})=u_{j}(e_{jk})$.
On the other hand, in the case depicted by Figure 2(c), the orientation $\pi$ directs the edge $e_{ij}$ toward $i$. So, $i$ sees no difference between $\pi$ and $\pi^{o}$ locally between $i$ and $j$, and $j$ thinks that it has given an edge of utility $u_{j}(e_{jk})$ to $i$ by going from $\pi^{o}$ to $\pi$.
It follows that for each pair of adjacent vertices $i,j$ of $G$, the vertex $i$, having possibly given some of its edges in $\pi^{o}$ to $j$ going from $\pi^{o}$ to $\pi$, does not strongly envy $j$ in $\pi$ because it does not strongly envy $j$ in $\pi^{o}$. Thus, no vertex of $G$ strongly envies any of its neighbours in $\pi$, so $\pi$ is EFX0. ∎
[…]
Figure 2: The three possibilities involving the fake edges $e_{ik},e_{jk}$ in the EFX0 orientation $\pi^{o}$ of $G^{o}$.
See 1
Proof.
We show FindEFXOrientation to be such an algorithm. If $G$ admits an EFX0 orientation, then $G^{o}$ does as well by Lemma 4(1), so the subroutine FindEFXOrientObj produces such an orientation $\pi^{o}$ of $G^{o}$ by Lemma 3. In this case, FindEFXOrientation outputs $\pi$, which is an EFX0 orientation of $G$ by Lemma 4(2). Otherwise, $G$ does not admit an EFX0 orientation, so so $G^{o}$ admits no EFX0 orientation by Lemma 4(2), causing the subroutine FindEFXOrientObj on Line 16 to output false and FindEFXOrientation to also output false.
We now analyze the running time. Initializing $(G^{o},u^{o})$, $O$, and $S$ on Lines 13 takes $O(|V(G)|^{2})$ time because we copy $G$ and subdivide each edge at most once. Calling FindEFXOrientObj on Line 16 takes $O(|V(G^{o})|^{2})$ time by Lemma 2(3). Since $|V(G^{o})|\leq 2|V(G)|$, we have $O(|V(G^{o})|^{2})=O(|V(G)|^{2})$. Finally, each of the final two for-loops on Lines 2021 and Lines 2226 takes $O(|E(G)|)$ time because $O\subseteq E(G)$ and $S\subseteq E(G)$. Thus, FindEFXOrientation runs in $O(|V(G)|^{2})$ time. ∎
[…]
Input: An instance $(G,u)$ of EFX0-Orientation.
Output: An EFX0 orientation $\pi$ of $G$ if one exists and false otherwise.
1:$(G^{o},u^{o})\leftarrow$ copy of $(G,u)$
2:$O\leftarrow\emptyset$
3:$S\leftarrow\emptyset$
4:for each edge $e_{ij}=\{i,j\}$ of $G$ do
5:     if exactly one of $u_{i}(e_{ij}),u_{j}(e_{ij})$ is negative then
6:         if $u_{i}(e_{ij})<0$ then
7:              $i\leftrightarrow j$. $\triangleright$ ensuring $u_{i}(e_{ij})=0$
8:         add new vertex $k$ to $G^{o}$
9:         add new edges $e_{ik}=\{i,k\},e_{jk}=\{j,k\}$ to $G^{o}$
10:         delete $e_{ij}$ from $G^{o}$
11:         $S\leftarrow S\cup\{(\{e_{ij}\},\{e_{ik}\})\}$
12:         $u^{o}_{i}(e_{ik}),u^{o}_{k}(e_{ik})\leftarrow u_{i}(e_{ij})$
13:         $u^{o}_{j}(e_{jk}),u^{o}_{k}(e_{jk})\leftarrow u_{j}(e_{ij})$
14:     else
15:         $O\leftarrow O\cup\{e\}$
16:$\pi^{o}\leftarrow\textsc{FindEFXOrientObj}(G^{o},u^{o})$triangleright Algorithm 2
17:if $\pi^{o}=\texttt{false}$ then
18:     return false
19:$\pi\leftarrow(\pi_{1},\pi_{2},\dots,\pi_{|V(G)|})$ where each $\pi_{i}=\emptyset$
20:for each edge $e_{ij}\in O$ do
21:     $\pi$ orients $e_{ij}$ the same way as $\pi^{o}$
22:for each ordered pair ($(\{e_{ij}\},\{e_{ik}\})\in S$ do
23:     if $\pi^{o}$ allocates $e_{ik}$ to $i$ then
24:         Let $\pi$ allocate $e_{ij}$ to $i$.
25:     else
26:         Let $\pi$ allocate $e_{ij}$ to $j$.
27:return $\pi$
4 EFX0 Orientations of Multigraphs
In this section, we turn our attention to multigraphs and consider the problem of deciding if a multigraph admits an EFX0 orientation, which we call EFX0-Orientation-Multigraphs. Our proof relies on a reduction from the NP-complete problem Partition KARP (1972).
See 2
Proof.
Let $S=\{s_{1},s_{2},\dots,s_{n}\}$ be an instance of Partition. Construct the multigraph $G$ on two vertices $a,b$ as follows. For each $s_{i}\in S$, we construct an edge $e_{i}$ between $a$ and $b$ and set $u_{a}(e_{i})=u_{b}(e_{i})=-s_{i}$. In addition, we create two self-loops $e_{a},e_{b}$ at each of $a,b$, respectively, and set $u_{a}(e_{a})=0$ and $u_{b}(e_{b})=0$. Finally, we assume that the utility functions $u_{a},u_{b}$ are additive.
We claim that there exists a partition $S=S_{a}\cup S_{b}$ such that $\sum_{x\in S_{a}}x=\sum_{x\in S_{b}}x$ if and only if $G$ admits an EFX0 orientation. Suppose $S=S_{a}\cup S_{b}$ is such a partition. Let $\pi=(\pi_{a},\pi_{b})$ be the orientation of $G$ in which $\pi_{a}=\{e_{a}\}\cup\{e_{i}:s_{i}\in S_{a}\}$ and $\pi_{b}=\{e_{b}\}\cup\{e_{i}:s_{i}\in S_{b}\}$. In other words, both self-loops $e_{a},e_{b}$ are directed toward their respective endpoints, and the edges between $a$ and $b$ are divided according to the partition $S=S_{a}\cup S_{b}$. Since $u_{a}$ and $u_{b}$ are additive, we have $u_{a}(\pi_{a})=-\sum_{x\in S_{a}}x=-\sum_{x\in S_{b}}x=u_{b}(\pi_{b})$, so $\pi$ is EFX0.
Conversely, suppose $G$ admits an EFX0 orientation $\pi=(\pi_{a},\pi_{b})$. Let $S_{a}$ (resp. $S_{b}$) denote the set of edges between $a,b$ that are directed toward $a$ (resp. $b$) in the orientation $\pi$. Clearly, $S_{a}\cup S_{b}$ is a partition of $S$. We claim that $\sum_{x\in S_{a}}x=\sum_{x\in S_{b}}x$. Because $\pi$ is an orientation and $e_{a},e_{b}$ are self-loops, we have $e_{a}\in\pi_{a}$ and $e_{b}\in\pi_{b}$. In particular, the definition of EFX0 implies $u_{a}(\pi_{a}\setminus\{e_{a}\})\leq u_{a}(\pi_{b})$. Moreover, since $e_{b}$ is a self-loop at $b$ and has zero utility to $a$, we have $u_{a}(\pi_{b})=u_{a}(\pi_{b}\setminus\{e_{b}\})$. Combining this with the previous inequality yields $u_{a}(\pi_{a}\setminus\{e_{v}\})\leq u_{a}(\pi_{b}\setminus\{e_{b}\})$. Observe that $\pi_{a}\setminus\{e_{a}\}$ is exactly $S_{a}$ and $\pi_{b}\setminus\{e_{b}\}$ is exactly $S_{b}$. Thus, $u_{a}(S_{a})\leq u_{a}(S_{b})$. The definition of the edges between $a$ and $b$ in $G$ implies $\sum_{x\in S_{a}}x=u_{a}(S_{a})\leq u_{a}(S_{b})=\sum_{x\in S_{b}}x$. On the other hand, the symmetry of $G$ implies that $\sum_{x\in S_{b}}x\leq\sum_{x\in S_{a}}x$ as well, so $\sum_{x\in S_{a}}x=\sum_{x\in S_{b}}x$. ∎
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), funding reference number RGPIN-2022-04518.