### v2

# Polynomial-Time Algorithms for Fair Orientations of Chores

Abstract
This paper addresses the problem of finding fair orientations of graphs of chores, in which each vertex corresponds to an agent, each edge corresponds to a chore, and a chore has zero marginal utility to an agent if its corresponding edge is not incident to the vertex corresponding to the agent. Recently, Zhou et al. (IJCAI, 2024) analyzed the complexity of deciding whether graphs containing a mixture of goods and chores have EFX orientations, and conjectured that deciding whether graphs containing only chores have EFX orientations is NP-complete. We resolve this conjecture by giving polynomial-time algorithms that find EF1 and EFX orientations of graphs containing only chores if they exist, even if there are self-loops. Remarkably, our result demonstrates a surprising separation between the case of goods and the case of chores, because deciding whether graphs containing only goods have EFX orientations was shown to be NP-complete by Christodoulou et al. (EC, 2023). In addition, we show the EF1 and EFX orientation problems for multigraphs to be NP-complete.
[…]
0561
1 Introduction
Consider the problem of dividing a set of items among a group of agents with the stipulation that each item must be given to exactly one agent and cannot be shared. This problem models a wide variety of settings, such as assigning offices in a building to employees, allocating classrooms to multiple simultaneous lectures, and even divorce settlements between spouses. The classical example of dividing a single good between two agents shows that it is not always possible to find an envy-free allocation, necessitating the search for alternative criteria of fairness.
The most prominent alternative is envy-freeness up to any good (EFX) [5, 9], which allows for each agent $i$ to envy each agent $j$, as long as the envy can be alleviated if the good in $j$’s bundle that has the smallest marginal utility for $i$ is removed from $j$’s bundle. The existence problem of EFX allocations has proven to be notoriously difficult and has been referred to as “fair division’s most enigmatic question” by Procaccia [15]. Fruitful results have been obtained by restricting instances. For example, EFX allocations exist if the agents have identical utility functions or identical ranking of the goods [14], if the agents have lexicographic utility functions [10], or if there are at most 3 agents [6, 2].
Another possibility is to limit the number of agents who are interested in each item. The assumption that each item provides non-zero marginal utility to at most two agents leads to the graphical instances introduced by Christodoulou et al. [7]. These instances can be represented by a graph $G$ wherein each vertex (resp. edge) corresponds to an agent (resp. good), and assuming that whenever an edge is not incident to a vertex, it provides zero marginal utility to the agent corresponding to that vertex. In this setting, it is natural to ask whether EFX orientations exist, i. e. whether it is possible to allocate each edge to one of its endpoints while satisfying the EFX condition. Unfortunately, it is NP-complete to decide whether EFX orientations exist for instances containing only goods [7].
For any result that holds for goods, one naturally ask what happens if chores are introduced. To this end, Zhou et al. [18] considered the EFX orientation problem in the mixed manna setting involving both goods and chores, and obtained a variety of complexity results for different variants of EFX in this setting. They conclude with a discussion on the special case of instances containing only chores, which can be used to model situations in which tasks are to be allocated to agents, but some tasks may not suitable to an agent because the agent is far away physically. For instance, a food delivery platform should avoid assigning orders to workers who are far away. Similarly, in a rental apartment where each floor has a set of communal chores, one should avoid assigning a chore to a tenant of a different floor.
Two variants of EFX for chores have been defined. Let $\pi$ be an allocation, $\pi_{i}$ denote the set of chores given to $i$, and $u_{i}$ denote the utility function of agent $i$. We say that $\pi$ is
  • EFX_ if for each pair of agents $i\neq j$ such that $i$ envies $j$ and each $e\in\pi_{i}$ such that $u_{i}(e)<0$, we have $u_{i}(\pi_{i}\setminus\{e\})\geq u_{i}(\pi_{j})$; and
  • EFX0 if for each pair of agents $i\neq j$ such that $i$ envies $j$ and each $e\in\pi_{i}$ such that $u_{i}(e)\leq 0$, we have $u_{i}(\pi_{i}\setminus\{e\})\geq u_{i}(\pi_{j})$.
The difference is that while EFX_ only requires the envy be alleviated when ignoring chores with strictly negative marginal utility, EFX0 requires that the envy be alleviated even if we ignore a chore with zero marginal utility. Clearly, EFX0 is a strictly stronger condition than EFX_.
Zhou et al. [18] found a polynomial-time algorithm that decides whether an EFX_ orientation exists, and conjectured that deciding whether an EFX0 orientation exists is NP-complete, similar to the case involving only goods. (This conjecture appears in the arXiv version of their paper [13] and not the conference version [18].)
1.1 Our Contribution
We study the complexity of finding EF1 and EFX0 orientations of chores on graphs and multigraphs that possibly contain self-loops. Importantly, we resolve the conjecture made by Zhou et al. [18] by giving a polynomial-time algorithm that decides the existence of an EFX0 orientation of chores and outputs one if it exists. Specifically,
  1. 1.
    There exists a $O(|V(G)|^{2})$-time algorithms that decide whether a graph $G$ of chores has an EF1 orientation and outputs one if it exists (discussion in Section 3).
  2. 2.
    There exists a $O(|V(G)|^{4})$-time algorithms that decide whether a graph $G$ of chores has an EFX0 orientation and outputs one if it exists (Theorem 9).
  3. 3.
    Deciding whether a multigraph of chores has an EF1 (resp. EFX0) orientation is NP-complete (Theorems 10 and 12).
Remarkably, our results for chores stand in contrast with previous results for goods. While EF1 orientations of goods always exist [8], our results imply they can fail to exist for chores. Second, while deciding whether EFX0 orientations of goods exist is NP-complete [7], out results imply the same problem is polynomial-time decidable for chores.
1.2 Related Work
While EFX allocations always exist [7], deciding whether EFX orientations exist is NP-complete [7], even for graphs with a vertex cover of size 8 [8]. EFX orientability has also been shown to be related to the chromatic number $\chi(G)$ of $G$ — while they always exist if $\chi(G)\leq 2$, they can fail to exist if $\chi(G)\geq 3$ [17].
The multigraph setting has also been studied. While deciding whether EFX orientations exist is NP-complete for multigraphs with only 10 vertices [8], some positive results have also been obtained. A multigraph is bivalued if there exists two values $\alpha$ and $\beta$ such that $u_{i}(e)\in\{\alpha,\beta\}$ for each vertex $i$ and each edge $e$. A multigraph is symmetric if for each edge $e=\{i,j\}$, we have $u_{i}(e)=u_{j}(e)$. EFX orientations exist for bivalued symmetric multigraphs not containing a structure called a non-trivial odd multitree [11]. EFX allocations of a multigraph $G$ exist if $G$ is bipartite [1, 4, 16], a multicycle [1], a multitree [4], if $G$ has girth $2\chi(G)-1$ [4], if each vertex has at most $\lceil|V(G)|/4\rceil-1$ neighbours [16], or if the shortest cycle with non-parallel edges has length at least 6 [16].
1.3 Organization
We introduce the model in Section 2. We give our algorithms for finding EF1 and EFX0 orientations of graphs in Sections 3 and 4. We prove our hardness results (Theorem 10 and Theorem 12) in Section 5. We discuss the special case of additive utility functions in Section 6. Finally, we conclude the paper in Section 7.
2 Preliminaries
2.1 The Fair Orientation Problem
An instance of the fair orientation problem is an ordered pair $(G,u)$ where $G=(V(G),E(G))$ is a graph and $u$ is a collection of $|V(G)|$ utility functions $u_{i}:\mathcal{P}(E(G))\to\mathbb{R}_{\leq 0}$. We allow for self-loops. Each vertex (resp. edge) corresponds to an agent (resp. chore). For each agent $i$, the function $u_{i}$ represents its preferences by assigning a real number to each subset of the edges. In the case of a single edge $e$, we write $u_{i}(e)$ in lieu of $u_{i}(\{e\})$ for brevity. We assume that whenever a vertex $i$ is not incident to an edge $e$, the edge $e$ has zero marginal utility to $i$, i. e. $u_{i}(S\cup\{e\})=u_{i}(S)$ for any set $S\subseteq E(G)$. In addition, we make the following two minimal assumptions on each utility function $u_{i}$:
  • $u_{i}$ is monotone, i. e., $u_{i}(S)\leq u_{i}(T)$ whenever $T\subseteq S$.
  • If $u_{i}(e)=0$ for each $e\in S$, then $u_{i}(S)=0$.
An allocation $\pi=(\pi_{1},\pi_{2},\dots,\pi_{|V(G)|})$ is an ordered collection of $|V(G)|$ pairwise disjoint subsets of $E(G)$ whose union is exactly $E(G)$. For each $i$, the subset $\pi_{i}$ represents the set of edges assigned to $i$. An allocation is called an orientation if there is a graph orientation $O$ of $G$ such that $e\in\pi_{i}$ if and only if $e$ is directed toward $i$ in $O$.
Definition 1 (EF1).
An allocation (resp. orientation) $\pi$ is EF1 if for each pair of agents $i\neq j$, there exists an edge $e\in\pi_{i}$ such that $u_{i}(\pi_{i}\setminus\{e\})\geq u_{i}(\pi_{j})$.
We say an agent $i$ strongly envies an agent $j$ if there exists an edge $e\in\pi_{i}$ such that $u_{i}(\pi_{i}\setminus\{e\})<u_{i}(\pi_{j})$.
Definition 2 (EFX0).
An allocation (resp. orientation) $\pi$ is EFX0 if no agent strongly envies another.
Given $(G,u)$, EFX0-Orientation asks whether $G$ has an EFX0 orientation. We are particularly interested in instances in which every edge $e=\{i,j\}$ is objective, i. e. $u_{i}(e)=0$ if and only if $u_{j}(e)=0$. Equivalently, either $e$ has zero marginal utility for both endpoints or $e$ has negative marginal utility for both. If every edge is objective, we call $(G,u)$ an objective instance of EFX0-Orientation. We use EFX0-Orientation-Objective to denote the EFX0-Orientation problem restricted to objective instances.
When considering objective instances, we refer to edges with zero (resp. negative) marginal utility for both endpoints as dummy edges (resp. negative edges). Moreover, we define a negative component of $G$ to be a maximal vertex-induced subgraph $K$ such that for any two vertices $i,j$ of $K$, there is a path in $K$ between $i$ and $j$ that only contains negative edges. The negative components of $G$ form a partition of $V(G)$. See Figure 1 for an example.
Figure 1: The graph of an objective instance. Solid (resp. dashed) edges denote negative (resp. dummy) edges. The set of black (resp. white) vertices induces one negative component.
2.2 Related Decision Problems
Let $H$ be a graph, $P=\{P_{1},P_{2},\dots,P_{k}\}$ be a set of pairwise disjoint subsets of $V(H)$, and $D\subseteq V(H)$ be a subset of vertices. The set $P$ is not necessarily a partition of $V(H)$ because not every vertex needs to be in $P$. We say a subset $C\subseteq V(H)$ is a $(P,D)$-vertex cover of $H$ if $C$ is a vertex cover of $H$ that contains at most one vertex from each $P_{\ell}\in P$ and no vertex in $D$, i. e. the following conditions hold:
  • Each edge of $H$ is incident to at least one vertex in $C$;
  • $|C\cap P_{\ell}|\leq 1$ for each $P_{\ell}\in P$; and
  • $C\cap D=\emptyset$.
Definition 3 (PD-Vertex-Cover).
Given a tuple $(H,P,D)$ where $H$ is a graph, $P=\{P_{1},P_{2},\dots,P_{k}\}$ is a set of pairwise disjoint subsets of $V(H)$, and $D\subseteq V(H)$, PD-Vertex-Cover asks whether there exists a $(P,D)$-vertex cover $C$ of $H$.
Definition 4 (2SAT).
The 2SAT problem is a special case of SAT in which each clause consists of one or two literals.
Definition 5 (Partition).
Given a set $S=\{s_{1},s_{2},\dots,s_{k}\}$ of natural numbers, Partition asks whether there exists a partition $S=S_{a}\cup S_{b}$ such that $\sum_{x\in S_{a}}x=\sum_{x\in S_{b}}x$. (Such a partition is called an equipartition.)
3 EF1 Orientations of Graphs
We give our $O(|V(G)|^{2})$-time algorithm for finding an EF1 orientation of a graph $G$ if one exists. Our algorithm follows immediately from the following two propositions. The second follows from the first and is stated without proof.
Proposition 1.
An orientation $\pi$ of a graph $G$ is EF1 if and only if each vertex $i$ receives at most one edge of negative utility to it.
Proof.
Suppose $i$ receives two edges $e=\{i,j\}$ and $e^{\prime}\neq e$ of negative utility to it. Assume $e^{\prime}$ is a chore in $\pi_{i}$ with maximum negative marginal utility to $i$. Since $e\in\pi_{i}\setminus\{e^{\prime}\}$, we have $u_{i}(\pi_{i}\setminus\{e^{\prime}\})\leq u_{i}(e)<0=u_{i}(\pi_{j})$, so $i$ envies $j$ even after ignoring $e^{\prime}$ and $\pi$ is not EF1. Otherwise, $i$ receives at most one edge of negative utility to it. Ignoring such an edge alleviates any envy $i$ has, so $\pi$ is EF1. ∎
Proposition 2.
If $\pi$ is an EF1 orientation of a graph $G$, then introducing a new edge $e=\{i,j\}$ such that $u_{i}(e)=0$ and orienting it toward $i$ results in another EF1 orientation. ∎
We also make use of the following graph theoretical observation.
Observation 3.
A graph $G$ has an orientation such that every vertex has in-degree at most one if and only if each component $K$ of a graph $G$ contains at most $|V(K)|$ edges. Moreover, such an orientation can be found in $O(|V(G)|^{2})$ time if it exists.
Proof.
Suppose each component $K$ of a graph $G$ contains at most $|V(K)|$ edges. Using breadth-first search (BFS), find a rooted spanning tree $T$ of each component $K$ of $G$ while orienting all tree edges away from the root $r(K)$ in its component. Each component $K$ contains at most one unoriented edge because $K$ contains at most $|V(K)|$ edges. For an unoriented edge $vw$ in a component $K$, orient $vw$ toward $w$ and reverse all edges on the directed path from the root $r(K)$ to $w$. The resulting orientation satisfies the above property in $O(|V(G)|^{2})$ time. Otherwise, some component $K$ contains >|V(K)|>|V(K)| edges and the pigeonhole principle prevents such an orientation from existing. ∎
To decide whether $G$ has an EF1 orientation and compute one if it exists, assume without loss of generality that every edge has strictly negative utility to both endpoints using Proposition 2. If each component $K$ of $G$ has at most $|V(K)|$ edges, then we can compute an orientation such that every vertex has in-degree at most one using Observation 3. Such an orientation is EF1 by Proposition 1. Otherwise, some component $K$ has >|V(K)|>|V(K)| edges, so Observation 3 implies that every orientation of $G$ contains a vertex with in-degree $\geq 2$. By Proposition 1, such orientations are not EF1. The above procedure clearly takes $O(|V(G)|^{2})$ time.
4 EFX0 Orientations of Graphs
In this section, we give our $O(|V(G)|^{4})$-time algorithm FindEFXOrientation (Algorithm 3) for EFX0-Orientation. Our algorithm is essentially a reduction from EFX0-Orientation to 2SAT. It comprises three nested routines, each functioning as a reduction from one decision problem to another. See Figure 2 for a visual overview.
We present the proofs of correctness of the algorithms in their nested order, starting with innermost one, FindPDVertexCover (Algorithm 1).
$EFX0-Orientation$-ObjectivePD-Vertex-Cover2SATFindEFXOrientation (Algorithm 3)FindEFXOrientationObj (Algorithm 2)FindPDVertexCover (Algorithm 1)
Figure 2: The reductions our algorithms represent. Each node on the left is a decision problem. A directed edge between nodes is a reduction and is labelled with the relevant algorithm.
4.1 FindPDVertexCover (Algorithm 1)
FindPDVertexCover (Algorithm 1) accepts an instance $(H,P,D)$ of PD-Vertex-Cover as input, and outputs a $(P,D)$-vertex cover of $H$ if one exists and false otherwise.
Recall that in an instance $(H,P,D)$ of PD-Vertex-Cover, $H$ is a graph, $P=\{P_{1},P_{2},\dots,P_{k}\}$ is a set of pairwise disjoint subsets of $V(H)$, and $D\subseteq V(H)$. The algorithm starts by constructing an instance $\phi$ of 2SAT on Line 1, which contains a Boolean variable $x_{i}$ for each vertex $i$ of $H$ and three types of clauses:
  • (type 1) the clause $\{x_{i},x_{j}\}$ for each edge $e=\{i,j\}$ of $H$ (possibly $i=j$),
  • (type 2) the clause $\{\overline{x_{i}},\overline{x_{j}}\}$ for each set $P_{\ell}\in P$ and each pair of distinct vertices $i,j\in P_{\ell}$, and
  • (type 3) the clause $\{\overline{x_{i}}\}$ for each vertex $i\in D$.
Then, the algorithm determines whether $\phi$ has a satisfying assignment $f$ on Line 2. If so, the algorithm outputs the $(P,D)$-vertex cover $C\coloneqq\{i\in V(H)\mid f(x_{i})=\texttt{true}\}$. Otherwise, the algorithm outputs false.
Algorithm 1 $\textsc{FindPDVertexCover}(H,P,D)$
Input: An instance $(H,P,D)$ of PD-Vertex-Cover.
Output: A $(P,D)$-vertex cover $C$ of $H$ if one exists and false otherwise.
[…]
3:if $f=\texttt{false}$ then $\triangleright$ $\phi$ has no satisfying assignment
4:  return false
5:return $C\coloneqq\{i\in V(H)\mid f(x_{i})=\texttt{true}\}$
Lemma 4.
The following statements hold for FindPDVertexCover (Algorithm 1):
  1. 1.
    If $H$ has a $(P,D)$-vertex cover, then $\phi$ has a satisfying assignment.
  2. 2.
    If $\phi$ has a satisfying assignment $f$, then $C\coloneqq\{i\in H\mid f(x_{i})={\texttt{true}}\}$ is a $(P,D)$-vertex cover of $H$. Otherwise, FindPDVertexCover outputs false.
  3. 3.
    FindPDVertexCover runs in $O(|V(H)|^{2})$ time.
Proof.
(1:) Suppose $H$ has a $(P,D)$-vertex cover $C$. We claim the truth assignment $f$ in which $f(x_{i})=\texttt{true}$ if and only if $i\in C$ to be a satisfying assignment for $\phi$. Since $C$ is a $(P,D)$-vertex cover of $H$, each edge $e$ of $H$ is incident to at least one vertex in $C$. In other words, for each edge $e=\{i,j\}$ of $H$, we have $i\in C$ or $j\in C$. Hence, at least one of $x_{i},x_{j}$ is true under $f$, so the type 1 clause $\{x_{i},x_{j}\}$ is satisfied.
On the other hand, the definition of $(P,D)$-vertex covers ensures $|C\cap P_{\ell}|\leq 1$ for each $P_{\ell}\in P$. Hence, each $P_{\ell}$ contains at most one vertex $i$ such that $f(x_{i})=\texttt{true}$, so all type 2 clauses are satisfied.
Finally, because $C$ is a $(P,D)$-vertex cover, we have $C\cap D=\emptyset$. Hence, for each $i\in D$, we have $i\notin C$, so $f(x_{i})=\texttt{false}$, ensuring all type 3 clauses are satisfied.
(2:) Suppose $\phi$ has a satisfying assignment $f$. We claim that $C\coloneqq\{i\in V(H)\mid f(x_{i})=\texttt{true}\}$ is a $(P,D)$-vertex cover of $H$. Since the type 1 clause $\{x_{i},x_{j}\}$ is satisfied under $f$, the edge $e=\{i,j\}$ of $H$ is incident to at least one vertex of $C$, namely, $i$ or $j$. Since the type 2 clauses are satisfied, for each fixed $P_{\ell}\in P$, there is at most one vertex $i\in P_{\ell}$ such that $x_{i}$ is true under $f$. Hence, $|C\cap P_{\ell}|\leq 1$. Finally, the type 3 clauses ensure that $f(x_{i})=\texttt{false}$ for each vertex $i\in D$, so $i\notin C$. Thus, $C\cap D=\emptyset$.
Otherwise, $\phi$ has no satisfying assignment and FindPDVertexCover outputs false on Line 4.
(3:) Constructing $\phi$ on Line 1 takes $O(|V(H)|^{2})$ time because $\phi$ contains $|V(H)|$ variables and $O(|V(H)|^{2})$ clauses. Verifying whether $\phi$ has a satisfying assignment on Line 2 takes time linear in the number of variables and clauses using the algorithm due to Aspvall et al. [3], which is $O(|V(H)|^{2})$. Finally, constructing $C$ takes $O(|V(H)|)$ time by checking each vertex of $H$. Thus, FindPDVertexCover runs in $O(|V(H)|^{2})$ time. ∎
4.2 FindEFXOrientObj (Algorithm 2)
FindEFXOrientObj (Algorithm 2) takes as input an instance $(G^{o},u^{o})$ of EFX0-Orientation-Objective and outputs an EFX0 orientation of $G^{o}$ if one exists and false otherwise. Recall that in an objective instance, all edges are objective, and we distinguish dummy edges that provide zero marginal utility to both endpoints and negative edges that provide negative marginal utility to both endpoints. Recall also that a negative component of $G^{o}$ is a maximal vertex-induced subgraph $K$ such that for any two vertices $i,j$ of $K$, there exists a path in $K$ between $i$ and $j$ that only contains negative edges.
We make the following crucial observation.
Proposition 5.
Let $(G^{o},u^{o})$ be an instance of EFX0-Orientation-Objective. An orientation $\pi^{o}$ of $G^{o}$ is EFX0 if and only for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge.
Proof.
Let $\pi^{o}=(\pi_{1}^{o},\pi_{2}^{o},\dots,\pi_{|V(G^{o})|}^{o})$ be an orientation of $G^{o}$ and suppose that for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. If $\pi^{o}$ contains a unique edge $e$ directed toward $i$, then we have $u_{i}^{o}(\pi^{o}_{i}\setminus\{e\})=u_{i}^{o}(\emptyset)=0\geq u_{i}^{o}(\pi_{j}^{o})$ for any agent $j$. Otherwise, every edge directed toward $i$ is a dummy edge, so $u_{i}^{o}(\pi_{i}^{o})=0\geq u_{i}^{o}(\pi_{j}^{o})$ for any $j$. Hence, no agent strongly envies another and $\pi^{o}$ is EFX0.
Conversely, suppose the orientation $\pi^{o}$ contains two edges $e,e^{\prime}$ directed toward some $i$ and not every edge directed toward $i$ is a dummy edge. Without loss of generality, assume $u_{i}^{o}(e^{\prime})<0$. Let $j,j^{\prime}$ denote the other endpoints of the edges $e,e^{\prime}$, respectively. Note that $j\neq j^{\prime}$ because $G$ does not contain parallel edges. Also, $u_{i}^{o}(\pi_{j}^{o})=0$ because $i$ receives the edge $e$ between $i$ and $j$. Since $u_{i}^{o}(e^{\prime})<0$ and $\{e^{\prime}\}\subseteq\pi_{i}^{o}\setminus\{e\}$, the monotonicity of $u_{i}^{o}$ implies $u_{i}^{o}(\pi_{i}^{o}\setminus\{e\})\leq u_{i}^{o}(e^{\prime})<0=u_{i}^{o}(\pi_{j}^{o})$. So, $\pi^{o}$ is not EFX0. ∎
Proposition 5 gives a graph theoretical condition that is equivalent to the EFX0 condition, which FindEFXOrientObj exploits. The algorithm first finds the set $\mathcal{K}$ of negative components of $G^{o}$ on Line 1. If some negative component $K$ contains more than $|V(K)|$ negative edges, the algorithm concludes that $G^{o}$ has no EFX0 orientation and outputs false. Otherwise, every negative component $K$ has at most $|V(K)|$ negative edges, and the algorithm constructs the instance $(H,P,D)$ of PD-Vertex-Cover on Line 4 defined as:
  • $H$ is the graph with the same vertex set as $G^{o}$ and the set of dummy edges of $G^{o}$;
  • $P=\{K\in\mathcal{K}\mid K\text{ has exactly }|V(K)|-1\text{ negative edges}\}$;
  • $D$ is the set of vertices that belong to a negative component $K\in\mathcal{K}$ that has exactly $|V(K)|$ negative edges.
Using FindPDVertexCover (Algorithm 1) on Line 5, the algorithm tries to find a $(P,D)$-vertex cover $C$ of $H$. If successful, it uses $C$ to construct the orientation $\pi^{o}$ of $G^{o}$ defined below and outputs it. Otherwise, it outputs false.
We give the construction of the orientation $\pi^{o}$ of $G^{o}$. Because $C$ is a $(P,D)$-vertex cover of $H$, each edge $e$ of $H$ (equivalently, each dummy edge of $G^{o}$) is incident to a vertex $i\in C$. Orient $e$ toward $i$, with arbitrary tie-breaking when two such $i$ exist. We then orient the negative edges of $G^{o}$ using Observation 3, which implies $G^{o}$ has an orientation such that every vertex has in-degree at most one (i. e. each vertex receives at most one negative edge) because every negative component $K$ of $G^{o}$ has at most $|V(K)|$ negative edges.
Algorithm 2 $\textsc{FindEFXOrientObj}(G^{o},u^{o})$
Input: An instance $(G^{o},u^{o})$ of EFX0-Orientation-Objective.
Output: An EFX0 orientation $\pi^{o}$ of $G^{o}$ if one exists and false otherwise.
[…]
2:if $\exists K\in\mathcal{K}\mid K$ contains >|V(K)|>|V(K)| negative edges then
3:  return false
[…]
6:if $C=\texttt{false}$ then
7:  return false
8:return the orientation $\pi^{o}$ of $G^{o}$ defined above
Lemma 6.
The following statements hold for FindEFXOrientObj (Algorithm 2):
  1. 1.
    If $G^{o}$ has an EFX0 orientation, then no negative component $K$ of $G^{o}$ contains >|V(K)|>|V(K)| negative edges and $H$ has a $(P,D)$-vertex cover.
  2. 2.
    If $C$ is a $(P,D)$-vertex cover of $H$, then $\pi^{o}$ is an EFX0 orientation of $G^{o}$.
  3. 3.
    FindEFXOrientObj runs in $O(|V(G^{o})|^{2})$ time.
Proof.
(1:) Suppose $G^{o}$ has an EFX0 orientation $\pi^{o}$. First, we show that no negative component $K$ of $G^{o}$ contains >|V(K)|>|V(K)| negative edges. Fix a negative component $K$. By Proposition 5, for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. In particular, this is true for each vertex in $K$, so $\pi^{o}$ contains at most $|V(K)|$ negative edges directed toward a vertex of $K$. On the other hand, every negative edge of $K$ is directed toward a vertex of $K$ because $\pi^{o}$ is an orientation. It follows that $K$ contains at most $|V(K)|$ negative edges.
We claim that the set $C$ of vertices $i$ of $G^{o}$ such that $\pi^{o}$ contains a dummy edge directed toward $i$ is a $(P,D)$-vertex cover of $H$. First we show that each edge of $H$ is incident to at least one vertex in $C$. Let $e$ be an edge of $H$. By the definition of $H$, the edge $e$ is a dummy edge of $G^{o}$. Let $i$ be the endpoint toward which $e$ is directed in the orientation $\pi^{o}$ of $G^{o}$. The definition of $C$ implies $i\in C$, so $e$ is incident to at least one vertex in $C$.
Next, we show that $|C\cap K_{\ell}|\leq 1$ for each $K_{\ell}\in P$. Fix any negative component $K_{\ell}$ of $G^{o}$. Since $K_{\ell}$ is a negative component, it contains at least $|K_{\ell}|-1$ negative edges. Since $\pi^{o}$ is EFX0, Proposition 5 implies that for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. Since $K_{\ell}$ contains at least $|K_{\ell}|-1$ negative edges and each vertex of $K$ can have at most one negative edge directed towards it, the pigeonhole principle implies that $K_{\ell}$ contains at most one vertex to which dummy edges are directed in $\pi^{o}$. In other words, at most one vertex of $K_{\ell}$ is in $C$, i. e., $|C\cap K_{\ell}|\leq 1$.
Finally, we show that $C\cap D=\emptyset$. Suppose $i\in D$. Because $i\in D$, the negative component $K$ to which $i$ belongs contains $|V(K)|$ negative edges. Proposition 5 implies that for each vertex $j$, the orientation $\pi^{o}$ contains a unique edge directed toward $j$ or every edge directed toward $j$ is a dummy edge. In particular, each vertex of $K$ has at most one negative edge directed toward it. By the pigeonhole principle, each vertex of $K$ has exactly one negative edge of $K$ directed toward it. In particular, this is true for $i$, so Proposition 5 implies no dummy edge is directed toward $i$. Thus, $i\notin C$. It follows that $C\cap D=\emptyset$ and $H$ indeed has a $(P,D)$-vertex cover.
(2:) Suppose the algorithm finds a $(P,D)$-vertex cover $C$ of $H$ on Line 5. First, we show that $\pi^{o}$ indeed orients every edge of $G^{o}$. Recall that $H$ is the graph with the same vertex set as $G^{o}$ and the set of dummy edges of $G^{o}$. Since $C$ is a $(P,D)$-vertex cover of $H$, each edge of $H$ (thereby each dummy edge of $G^{o}$) is incident to at least one vertex in $C$. It follows that every dummy edge of $G^{o}$ is oriented by $\pi^{o}$. On the other hand, let $e$ be a negative edge of $G^{o}$ and $K$ be the negative component containing $e$. Since the algorithm successfully finds a $(P,D)$-vertex cover $C$ of $H$ on Line 5, it does not terminate on Line 3. So, $K$ contains exactly $|V(K)|$ or |V(K)|1|V(K)|-1 negative edges, and $\pi^{o}$ orients all of them. It follows that $\pi^{o}$ indeed orients every edge of $G^{o}$.
It remains to show $\pi^{o}$ to be EFX0. We verify the condition in Proposition 5 for every vertex $i$. Let $C^{\prime}\subseteq C$ be the set of vertices with a dummy edge directed toward it in $\pi^{o}$ and consider any $i\in C^{\prime}$. Since $C$ is a $(P,D)$-vertex cover of $H$, we have $C\cap D=\emptyset$, so $i\notin D$. Because $D$ is the set of vertices that belong to a negative component $K$ of $G^{o}$ that has exactly $|V(K)|$ negative edges, $i$ belongs to a negative component $K$ of $G^{o}$ containing exactly |V(K)|1|V(K)|-1 negative edges. As discussed while defining $\pi^{o}$, at most one vertex of such a negative component $K$ could have received a dummy edge, which clearly is the vertex $i$. So, $\pi^{o}$ orients the negative edges of $K$ in the direction away from $i$, so that they form a spanning tree of $K$ rooted at $i$. Hence, every edge directed toward $i$ is a dummy edge. it follows that the condition in Proposition 5 holds for every $i\in C^{\prime}$.
On the other hand, the definition of $\pi^{o}$ ensures each vertex receives at most one negative edge. In particular, each vertex $i\notin C^{\prime}$ receives no dummy edges and at most one negative edge, so the condition in Proposition 5 holds for every $i\notin C^{\prime}$ as well. By Proposition 5, $\pi^{o}$ is EFX0.
(3:) Finding the set $\mathcal{K}$ of negative components of $G^{o}$ on Line 1 can be done using a BFS using $O(|V(G^{o})|+|E(G^{o})|)$ time. Constructing $(H,P,D)$ on Lines 4 requires counting the number of negative edges in each negative component $K\in\mathcal{K}$, which can be done in $O(|E(G^{o})|)$ time. By Lemma 4(3), FindPDVertexCover on Line 5 runs in $O(|V(H)|^{2})$ time.
The construction of $\pi^{o}$ on Line 8 is done in two steps. First, orienting the dummy edges can be done in $O(|V(G^{o})|+|E(G^{o})|)$ time because there it requires checking which endpoint of each dummy edge is in the $(P,D)$-vertex cover $C$. Then, orienting the negative edges can be done in $O(|V(G^{o})|^{2})$ time according to Observation 3.
By summing the above and observing that $|V(H)|=|V(G^{o})|$, we conclude that FindEFXOrientObj runs in $O(|V(G^{o})|^{2})$ time. ∎
Lemma 7.
Given an instance $(G^{o},u^{o})$ of EFX0-Orientation-Objective, FindEFXOrientObj outputs an EFX0 orientation $\pi^{o}$ of $G^{o}$ if one exists and false otherwise in $O(|V(G^{o})|^{2})$ time.
Proof.
If $G^{o}$ has an EFX0 orientation, then no negative component $K$ of $G^{o}$ contains >|V(K)|>|V(K)| negative edges and $H$ has a $(P,D)$-vertex cover by Lemma 6(1). So, Line 5 finds a $(P,D)$-vertex cover $C$ of $H$ by Lemma 4 and FindEFXOrientObj outputs the EFX0 orientation $\pi^{o}$ by Lemma 6(2).
Otherwise, $G^{o}$ does not have an EFX0 orientation. If some negative component $K$ of $G^{o}$ contains >|V(K)|>|V(K)| negative edges, then FindEFXOrientObj outpts false on Line 3, Assume every negative component $K$ of $G^{o}$ contains at most $|V(K)|$ negative edges, so FindEFXOrientObj constructs the instance $(H,P,D)$ of PD-Vertex-Cover on Line 4 and attempts to find a $(P,D)$-vertex cover $C$ of $H$ on Line 5. By the contrapositive of Lemma 6(2), such a $C$ does not exist, so FindEFXOrientObj outputs false on Line 7. ∎
4.3 FindEFXOrientation (Algorithm 3)
FindEFXOrientation (Algorithm 3) accepts an instance $(G,u)$ of EFX0-Orientation as input and outputs an EFX0 orientation of $G$ if one exists and false otherwise.
It first constructs the instance $(G^{o},u^{o})$ of EFX0-Orientation-Objective by subdividing each non-objective edge of $G$. Specifically, for each non-objective edge $e_{ij}=\{i,j\}$, first assume $u_{i}(e_{ij})=0$ without loss of generality by possibly exchanging the labels of $i$ and $j$, then replace $e_{ij}$ by a new vertex $k$ and two new edges $e_{ik}=\{i,k\},e_{jk}=\{j,k\}$. The utilities of the new edges are defined as shown in Figure 3, in which $\beta\coloneqq u_{j}(e_{ij})$. The objective edges and their utilities remain unchanged.
It then uses FindEFXOrientObj (Algorithm 2) on Line 2 to try to find an EFX0 orientation of $G^{o}$. If none exists, it outputs false on Line 4. Otherwise, an EFX0 orientation $\pi^{o}$ of $G^{o}$ is successfully found. Using $\pi^{o}$, the algorithm constructs and outputs the orientation $\pi$ of $G$ defined as follows.
Let $e_{ij}$ be an edge of $G$. If $e_{ij}$ was not subdivided in the construction of $G^{o}$, then it is also an edge in $G^{o}$. In this case, $\pi$ orients $e_{ij}$ in $G$ in the same way as $\pi^{o}$ orients $e_{ij}$ in $G^{o}$.
Otherwise, $e_{ij}$ was subdivided in the construction of $G^{o}$, and corresponds to two edges $e_{ik},e_{jk}$ of $G^{o}$, where $u_{i}^{o}(e_{ik})=u_{i}(e_{ij})=0$. In this case, $\pi$ orients $e_{ij}$ in $G$ in the same direction as $\pi^{o}$ orients $e_{ik}$ in $G^{o}$ (with respect to Figure 3). That is, if $\pi^{o}$ orients $e_{ik}$ toward $i$ in $G^{o}$, then $\pi$ orients $e_{ij}$ toward $i$ in $G$. Otherwise, $\pi^{o}$ orients $e_{ik}$ toward $k$ in $G^{o}$, and $\pi$ orients $e_{ij}$ toward $j$ in $G$.
[…]
(a) the edge $e_{ij}$ in $G$
[…]
(b) after subdivision in $G^{o}$
Figure 3: Subdivision of $e_{ij}$ during the construction of $G^{o}$ by FindEFXOrientation. The labels above an edge indicate the utility the edge has to its two endpoints. We write $\beta\coloneqq u_{j}(e_{ij})$ for clarity.
Algorithm 3 $\textsc{FindEFXOrientation}(G,u)$
Input: An instance $(G,u)$ of EFX0-Orientation.
Output: An EFX0 orientation $\pi$ of $G$ if one exists and false otherwise.
[…]
3:if $\pi^{o}=\texttt{false}$ then
4:  return false
5:return $\pi\coloneqq$ the orientation of $G$ defined above
Lemma 8.
The following statements hold for FindEFXOrientation (Algorithm 3):
  1. 1.
    If $G$ has an EFX0 orientation, then so does $G^{o}$.
  2. 2.
    If $\pi^{o}$ is an EFX0 orientation of $G^{o}$, then $\pi$ is an EFX0 orientation of $G$.
Proof.
(1:) Suppose $\pi$ is an EFX0 orientation of $G$. We construct the orientation $\pi^{o}$ of $G^{o}$ as follows. For each edge $e$ of $G$ that was not subdivided during the construction of $G^{o}$, we let $\pi^{o}$ orient it in $G^{o}$ in the same direction that $\pi$ orients it in $G$. On the other hand, suppose $e_{ik}$ and $e_{jk}$ are a pair of edges of $G^{o}$ that resulted from the subdivision of the edge $e_{ij}$ of $G$. If $\pi$ orients $e_{ij}$ leftward in terms of Figure 3 (resp. rightward), then $\pi^{o}$ orients both $e_{ik}$ and $e_{jk}$ leftward (resp. rightward).
We claim $\pi^{o}$ is EFX0. We refer to a vertex that is shared between both $G$ and $G^{o}$ as a real vertex, and a vertex that exists only in $G^{o}$ as a fake vertex. Fake vertices result from the subdivisions during the construction of $G^{o}$. Clearly, from the local perspective of each real vertex, the orientations $\pi$ and $\pi^{o}$ appear the same. Since no real vertex strongly envies a neighbour in $\pi$, no real vertex strongly envies a neighbour in $\pi^{o}$. On the other hand, the construction of $\pi^{o}$ ensures each fake vertex receives exactly one incident edge, so no fake vertex strongly envies a neighbour in $\pi^{o}$. Thus, $\pi^{o}$ is EFX0.
(2:) Suppose $\pi^{o}$ is an EFX0 orientation of $G^{o}$. Proposition 5 implies for each vertex $i$, the orientation $\pi^{o}$ contains a unique edge directed toward $i$ or every edge directed toward $i$ is a dummy edge. In particular, for any fake vertex $k$ in $G^{o}$ and the two edges $e_{ik},e_{jk}$ incident to it, $e_{ik}$ is a dummy edge and $e_{jk}$ is a negative edge, so at most one of them is directed toward $k$ in $\pi^{o}$. Hence, there are only three possibilities for how pairs of fake edges are oriented in $\pi^{o}$ (see Figure 4).
Consider the first two cases depicted in Figure 4. In Figure 4(a), both $e_{ik}$ and $e_{jk}$ are directed rightward in $\pi^{o}$, and $\pi$ orients the corresponding edge $e_{ij}$ rightward. In Figure 4(b), both $e_{ik}$ and $e_{jk}$ are directed leftward in $\pi^{o}$, and $\pi$ orients $\pi_{ij}$ leftward. In either case, the real vertices $i$ and $j$ see no difference between $\pi$ and $\pi^{o}$ locally in terms of the utilities they derive from the edges $e_{ij},e_{ik},e_{jk}$ between $i$ and $j$ that they receive, because $u_{i}(e_{ij})=u_{i}(e_{ik})$ and $u_{j}(e_{ij})=u_{j}(e_{jk})$.
In the case depicted by Figure 4(c), the orientation $\pi$ directs the edge $e_{ij}$ toward $i$. So, $i$ sees no difference between $\pi$ and $\pi^{o}$ locally between $i$ and $j$, and $j$ thinks that it has given an edge of utility $u_{j}(e_{jk})$ to $i$ by going from $\pi^{o}$ to $\pi$.
Hence, for each pair of adjacent vertices $i,j$ of $G$, the vertex $i$, having possibly given some of its edges in $\pi^{o}$ to $j$ going from $\pi^{o}$ to $\pi$, does not strongly envy $j$ in $\pi$ because it does not strongly envy $j$ in $\pi^{o}$. Thus, $\pi$ is EFX0. ∎
[…]
Figure 4: The three possibilities involving the fake edges $e_{ik},e_{jk}$ in the EFX0 orientation $\pi^{o}$ of $G^{o}$.
Theorem 9.
There exists a $O(|V(G)|^{4})$-time algorithm that decides whether a graph $G$ of chores has an EFX0 orientation and outputs one if it exists.
Proof.
We show FindEFXOrientation to be such an algorithm. If $G$ has an EFX0 orientation, then $G^{o}$ does as well by Lemma 8(1), so the subroutine FindEFXOrientObj produces such an orientation $\pi^{o}$ of $G^{o}$ by Lemma 7. In this case, FindEFXOrientation outputs $\pi$, which is an EFX0 orientation of $G$ by Lemma 8(2). Otherwise, $G$ does not have an EFX0 orientation, so so $G^{o}$ has no EFX0 orientation by Lemma 8(2), causing the subroutine FindEFXOrientObj on Line 2 to output false and FindEFXOrientation to also output false.
We now analyze the running time. Constructing $(G^{o},u^{o})$ takes $O(|V(G)|^{2})$ time because we subdivide each edge at most once. FindEFXOrientObj on Line 2 takes $O(|V(G^{o})|^{2})$ time by Lemma 6(3). Since $G^{o}$ is constructed by subdividing each edge of $G$ at most once, and each subdivision introduces a new vertex, we have $|V(G^{o})|\leq|V(G)|+|E(G)|=O(|V(G)|^{2})$. Hence, $O(|V(G^{o})|^{2})=O(|V(G)|^{4})$. Finally, constructing $\pi$ using $\pi^{o}$ takes $O(|E(G)|)$ time because it requires orienting each edge of $G$. Thus, FindEFXOrientation runs in $O(|V(G)|^{4})$ time. ∎
5 EF1 and EFX0 Orientations of Multigraphs
We turn our attention to multigraphs and consider the problem of deciding if a multigraph has EF1 or EFX0 orientations. We show that both of these problems are NP-complete using a reduction from the NP-complete problem Partition [12].
We present two different reductions. The first reduction (Theorem 10) is relatively simple but results in a multigraph with two self-loops. The second reduction (Theorem 12) is more complex but has the advantage of not using self-loops.
Theorem 10.
Deciding whether a multigraph $G$ of chores has an EF1 (resp. EFX0) orientation is NP-complete, even if $G$ is symmetric, has only two vertices, and utility functions are additive.
Proof.
Both of these problems are in NP because verifying whether a given orientation of a multigraph of chores is EF1 or EFX0 can clearly be done in polynomial time.
We reduce an instance $S=\{s_{1},s_{2},\dots,s_{k}\}$ of Partition to each of the EF1 and EFX0 orientation problems. We give the EF1 reduction and show how to adapt it for EFX0.
Let $\alpha<-\max_{i}s_{i}<0$. Construct a multigraph $G$ on two vertices $a,b$ as follows. For each $s_{i}\in S$, create an edge $e_{i}$ between $a$ and $b$ and set $u_{a}(e_{i})=u_{b}(e_{i})=-s_{i}$. Create two self-loops $e_{a},e_{b}$ at each of $a,b$, respectively, and set $u_{a}(e_{a})=u_{b}(e_{b})=\alpha$. Assume the utility functions are additive.
Partitions $S_{a}\cup S_{b}$ of $S$ correspond to orientations of the edges between $a$ and $b$ in a natural way. Moreover, because both $a$ and $b$ receive their respective self-loop which has $\alpha<-\max_{i}s_{i}<0$ utility, the EF1 criteria requires that their envy for each other be alleviated when they ignore their respective self-loops. Thus, $\sum_{x\in S_{a}}x=\sum_{x\in S_{b}}x$ if and only if $G$ has an EF1 orientation.
To adapt the reduction for EFX0, let $\alpha=0$ instead. ∎
We now show that deciding whether EF1 orientations exist remain NP-complete even if the multigraph $G$ contains no self-loops. To do this, we again rely on a reduction from Partition. For any instance $S=\{s_{1},s_{2},\dots,s_{k}\}$ of Partition, we define a multigraph $G$ on three vertices $a,b,c$ as follows. All of the edges $G$ contains has equal utility to both endpoints (called its weight). For each $s_{i}\in S$, create an edge between $a$ and $b$ of weight $s_{i}$. Then, between $c$ and each of $a,b$, create two edges of weight $T\coloneqq\sum_{i\in[k]}s_{i}$. (See Figure 5.)
Lemma 11.
In any EF1 orientation of $G$, at least one edge between $c$ and $a$ is oriented toward $a$ (and similarly between $c$ and $b$ by symmetry).
Proof.
Let $\pi=(\pi_{a},\pi_{b},\pi_{c})$ be an EF1 orientation of $G$. If both edges between $c$ and $a$ are oriented toward $c$, then for either one of them (call it $e$), we have $u_{c}(\pi_{c}\setminus\{e\})\leq-T<0=u_{c}(\pi_{a})$, so $\pi$ is not EF1. ∎
Theorem 12.
Deciding whether a multigraph $G$ of chores has an EF1 orientation is NP-complete, even if $G$ is symmetric, has only three vertices and no self-loops, and utility functions are additive.
Proof.
We have already seen in the proof of Theorem 12 that this problem is in NP. We now show that $S$ has an equipartition if and only if $G$ has an EF1 orientation. Suppose $S=S_{1}\cup S_{2}$ is an equipartition. Let $\pi$ be the orientation that orients the edges between $a$ and $b$ in the way that naturally corresponds to $S_{1}\cup S_{2}$, the two edges between $c$ and $a$ in opposite directions (one toward $c$ and the other toward $a$), and the two edges between $c$ and $b$ in opposite directions. We claim $\pi$ is EF1.
Consider the envy that $c$ experiences. It suffices to consider the envy from $c$ to $a$ by symmetry. By construction, $u_{c}(\pi_{c})=-2T<-T=u_{c}(\pi_{a})$. Because both edges directed toward $c$ have utility T-T, disregarding either of them alleviates the envy from $c$ to $a$.
Next, consider the envy that $a$ and $b$ experience. Again by symmetry, it suffices to consider only $a$. The edge $e$ between $a$ and $c$ directed toward $a$ has the greatest negative utility among the edges in $\pi_{a}$ by the definition of $T$, so the definition of EF1 requires the envy $a$ experiences to be alleviated if $e$ is disregarded. Indeed, we have $u_{a}(\pi_{a}\setminus\{e\})=(-1/2)T=u_{a}(\pi_{b})$ and $u_{a}(\pi_{a}\setminus\{e\})=(-1/2)T<-T=u_{a}(\pi_{c})$. So, $\pi$ is EF1.
Conversely, suppose $S$ has no equipartition. Let $\pi$ be any orientation of $G$ and $S_{1}\cup S_{2}$ be the partition of $S$ that corresponds to the way $\pi$ orients the edges between $a$ and $b$. More specifically, let $s_{i}\in S_{1}$ if and only if the edge between $a$ and $b$ corresponding to $s_{i}$ is directed toward $a$, and let $S_{2}=S\setminus S_{1}$.
If both edges between $c$ and $a$ are directed toward $c$ or if both edges between $c$ and $b$ are directed toward $b$, then $\pi$ is not EF1 by Lemma 11, so we are done. It remains to consider the case that an edge $e_{ac}$ between $c$ and $a$ is directed toward $a$ and an edge $e_{bc}$ between $c$ and $b$ is directed toward $b$. The definition of EF1 requires $\sum_{x\in S_{1}}x=u_{a}(\pi_{a}\setminus\{e_{ac}\})\leq u_{a}(\pi_{b})=-\sum_{x\in S_{2}}x$. Similarly, $\sum_{x\in S_{2}}x=u_{b}(\pi_{b}\setminus\{e_{bc}\})\leq u_{b}(\pi_{a})=-\sum_{x\in S_{1}}x$. So, $\sum_{x\in S_{1}}=\sum_{x\in S_{1}}x$, a contradiction as $S$ has no equipartition. ∎
[…]
Figure 5: The multigraph $G$ in Theorem 12. Here, $T\coloneqq\sum_{i\in[k]}s_{i}$.
6 On the Additive Case
All of our results also apply to the additive case in which every agent has an additive utility function, without any modification to the theorem statements and proofs. More generally, our results hold even if we allow each agent to have either a monotone utility function as previously described in Section 2 or an additive utility function, independently of other agents. We explain the reason below.
Our results pertaining to EF1 orientations stem from Proposition 1, which translates the EF1 orientation condition for each agent into a logically equivalent graph theoretical condition. Thus, any role that assumptions made on utility functions play lie entire within the proof of Proposition 1. It is easy to verify the correctness of this proof for additive utility functions as well.
As for EFX0 orientations, Proposition 5 translates the EFX0 orientation condition for each agent in an objective instances into an equivalent graph theoretical condition. So, we need to verify Proposition 5 and also the reduction of EFX0-Orientation to EFX0-Orientation-Objective for additive utility functions (i. e. Lemma 8). Again, it is straightforward to verify the proofs of both of these results for additive utility functions.
7 Conclusion
In this paper, we determined the complexity of finding EF1 and EFX0 orientations of chores in both the graph and multigraph settings, and gave polynomial-time algorithms for the polynomial-time cases.
{ack}
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), funding reference number RGPIN-2022-04518. We also thank the anonymous reviewers for their constructive comments.